# Adaptive Uncertainty Quantification for Generative AI

### v1

@@ C.1 Conformal Prediction for Regression
In this section, we detail the experimental procedures used in the creation of Table 1. For the dataset, besides the synthetic datasets (Data 1 and Data 2) in Section 4.1, we use (bike, bio, community, concrete, homes, star) for regression problems that have been previously used as benchmarks in conformal prediction [29].
C.1.1 Implementation Details
For all methods, we use a desired threshold of $\alpha=0.1$. We split each dataset randomly such that 40% of the data is used for training, 40% for calibration, and 20% for evaluation. The results in the table are means over ten trials with different splits.
For Conformal Tree, we set a maximum number of leaf nodes as $K=80$, a minimum number of samples per leaf as $m=50$, and a minimum improvement threshold of 0.05. We did not tune these hyperparameters, and used the same values for all datasets. Conformal Tree uses the robust tree algorithm (Algorithm 2) for partitioning, which satisfies the theoretical guarantees in the text. Conformal Tree (CART) uses standard CART for splitting, using the same hyperparemeters. Conformal Forest combines intervals from 100 CART trees, using bagging and random feature selection. The intervals are combined using a recent majority vote procedure referenced in the text [12]. Each tree in the forest also is regularized using the same hyperparameters as described earlier.
As the predictive model for all datasets, we used a random forest model with 100 estimators, fit using the variance reduction criterion, requiring a minimum of two samples to split. For CQR, the quantiles are predicted in the manner of Meinshausen and Ridgeway [21]. All other methods only utilize the predictions of the conditional mean from the model. Locally weighted (Forest) uses an auxiliary random forest model with the same hyperparameters predicting the absolute residuals of the original model on the training data.
C.2 Large Language Model Uncertainty Quantification
In these experiments, we use a large language model for downstream classification tasks. The first task is predicting a skin-disease from a description of physical symptoms. The second is predicting U.S. legislators’ state of employment from their DW-NOMINATE scores [24]. These are estimates of two-dimensional latent variables in a spatial model based on roll-call voting. To an unfamiliar reader, they can loosely interpreted as two-dimensional summary statistics pertaining to the legislators’ ideology.
C.2.1 Data
@@
We used the GPT-4o model produced by OpenAI in June of 2024 for our classification problems. We used a temperature setting of 0.0, yielding (ideally) deterministic output, except for our naive uncertainty quantification comparisons, in which we used a temperature setting of 1.0.
@@
The calibration data for the skin disease task is found in the UCI machine learning repository [14]. The dataset consists of 366 observations. We use 80% of the data for calibration and 20% for evaluation. The data consists of 35 features. For our experiments, we use the twelve clinical features, and ignore the remaining 23 histopathological features, as the purpose of this experiment is to emulate a user describing physical symptoms to a language model. The prompt used for the classification task is given below.
@@ 10 Format your answer as:
C.2.2 Conformal Prediction Details
For the skin disease example, we fixed the desired confidence level $\alpha$ at 0.1. We set a maximum number of leaves at $K=20$ and a minimum number of samples per leaf at $m=10$. We prefer a smaller value here to promote more local variation.
For the ideology example, we fix the desired confidence level $\alpha$ at 0.2, as the task is significantly more difficult. We set a maximum number of leaves at $K=20$ and a minimum number of samples per leaf at $m=200$. These are higher to represent the limited expected gain from partitioning the two-dimensional DW-NOMINATE space as opposed to the twelve-dimensional covariate space in the other example.
@@ C.2.2 Conformal Prediction Details
C.3 Additional Visualizations
In this section, we include additional visualizations referenced from the text. We include a plot of the empirical test coverage of split conformal and Conformal Tree on the GPT diagnosis example, as well another plot visualizing Conformal Tree’s tree structure, but with the empirical distributions of heldout test data below each node.
@@ C.3 Additional Visualizations
Figure 8: Empirical test coverage for split-conformal and Conformal Tree for varying $1-\alpha$ values on the GPT diagnosis example. Conformal tree always empirically covers at the desired level.
@@
Figure 9: First five levels of Conformal Tree’s structure on the skin disease test data. Node colors correspond to the mid-range of conformity scores (of calibration data) within the node. Below each node, we plot the empirical distribution of test data corresponding to each disease within that node. See Figure 7 for an analogous plot visualizing the calibration data.

### v2

@@ Appendix D Experimental Details
For the dataset, besides the synthetic datasets (Data 1 and Data 2) in Section 4.1, we use (bike, bio, community, concrete, homes, star) for regression problems that have been previously used as benchmarks in conformal prediction [44]. As the predictive model for all datasets, we used a random forest model with 100 trees, fit using the variance reduction criterion, requiring a minimum of two samples to split.
For Conformal Tree, we used a fixed minimum improvement threshold of 0.05, and tuned the hyperparameters $m$ and $K_{\mathrm{max}}$ on each dataset in order to minimize the interval score loss. Conformal Forest mentioned in Remark 3 combines intervals from 100 CART trees, with randomness arising from bagging and feature subsampling. The intervals are combined using a recent majority vote procedure referenced in the text [17]. Each tree in the forest also is regularized using the same hyperparameters as described earlier. The Locally weighted method uses an auxiliary random forest model to predict the absolute residuals. The hyperparameters of this model are tuned analogously, as we select the combination that minimizes interval score length on held-out validation data. For CQR, the quantiles are predicted in the manner of Meinshausen and Ridgeway [34]. All methods other than CQR only utilize the predictions of the conditional mean from the model.
D.2 Large Language Model Uncertainty Quantification
@@
In the experiments in Secion 5, we use a large language model for downstream classification tasks. We used the GPT-4o model produced by OpenAI in June of 2024 for our classification problems. We used a temperature setting of 0.0, yielding (ideally) deterministic output, except for our naive uncertainty quantification comparisons, in which we used a temperature setting of 1.0.
D.2.1 Skin-disease Prediction
@@
The calibration data for the skin disease task is found in the UCI machine learning repository [22]. The dataset consists of 366 observations. We use the twelve clinical features, and ignore the remaining 23 histopathological features, as the purpose of this experiment is to emulate a user describing physical symptoms to a language model. The prompt used for the classification task is given below.
@@ 10 Format your answer as:
D.2.2 Classifying States of Legislators
The task in this experiment is to predict U.S. legislators’ state of employment from their DW-NOMINATE scores [39]. These are estimates of two-dimensional latent variables in a spatial model based on roll-call voting. To an unfamiliar reader, they can loosely interpreted as two-dimensional summary statistics pertaining to the legislators’ ideology.
@@ D.2.2 Classifying States of Legislators
Appendix E Additional Experimental Results
In this section, we include additional visualizations referenced from the main body.
Constant calibration set size
We repeat the experiment in Section 4.2 with a fixed size of $n=500$ points in the calibration set for each approach, rather than a proportion of the total data size. The results are in Table 5 when a single tree is utilized and in Table 5 when a forest is used.
Split Conformal Conformal Tree Locally Weighted (Tree) CQR (Tree)
Width Cov. ISL Width Cov. ISL P.B. Width Cov. ISL P.B. Width Cov. ISL P.B.
Data 1 4.29 0.93 5.37 3.59 0.91 4.33 0.59 3.85 0.92 4.39 0.59 4.57 0.88 6.16 0.8
Data 2 2.69 0.9 5.07 2.87 0.87 3.85 0.87 3.31 0.91 4.27 0.58 5.29 0.9 9.51 0.0
bike 1.18 0.91 1.83 1.11 0.87 1.83 0.59 1.13 0.89 1.61 0.59 3.63 0.89 5.0 0.0
bio 1.53 0.89 2.12 1.48 0.88 2.15 0.48 1.62 0.91 2.14 0.29 3.41 0.91 3.61 0.0
community 1.83 0.9 2.77 1.68 0.88 2.6 0.67 1.82 0.91 2.47 0.59 3.03 0.89 4.66 0.0
concrete 0.67 0.91 0.87 0.63 0.88 0.88 0.62 0.65 0.9 0.89 0.53 1.97 0.88 2.37 0.0
homes 0.58 0.9 1.13 0.58 0.88 1.08 0.62 0.63 0.88 1.06 0.59 3.49 0.91 4.19 0.0
star 0.17 0.89 0.21 0.16 0.87 0.22 0.55 0.17 0.89 0.21 0.4 0.24 0.89 0.29 0.0
Table 4: Empirical results for tree-based methods on real and synthetic data as the average over five trials on width, coverage, ISL, and proportion better (P.B.) for $\alpha=0.1$. Calibration sets all have the size $n=500$.
Split Conformal Conformal Forest Locally Weighted (Forest) CQR (Forest)
Width Cov. ISL Width Cov. ISL P.B. Width Cov. ISL P.B. Width Cov. ISL P.B.
Data 1 4.13 0.9 5.84 3.64 0.9 4.34 0.62 3.77 0.9 4.56 0.62 3.13 0.86 4.38 0.7
Data 2 2.53 0.89 5.27 2.83 0.89 3.69 0.76 2.9 0.88 4.44 0.74 3.72 0.89 4.88 0.76
bike 1.23 0.91 1.81 1.19 0.94 1.43 0.54 1.08 0.91 1.64 0.67 2.1 0.89 2.56 0.26
bio 1.52 0.89 2.14 1.51 0.9 2.11 0.5 1.54 0.9 2.14 0.46 2.15 0.89 2.4 0.0
community 1.86 0.89 2.88 1.69 0.92 2.41 0.63 1.59 0.9 2.43 0.64 2.0 0.89 2.5 0.51
concrete 0.62 0.9 0.92 0.63 0.92 0.81 0.56 0.56 0.87 0.89 0.71 1.17 0.87 1.43 0.0
homes 0.62 0.91 1.15 0.66 0.93 0.94 0.64 0.56 0.91 1.02 0.68 1.19 0.88 1.93 0.26
star 0.17 0.89 0.21 0.17 0.9 0.21 0.36 0.17 0.89 0.21 0.51 0.19 0.91 0.22 0.18
Table 5: Empirical results for forest-based methods on real and synthetic data as the average over five trials on width, coverage, ISL, and proportion better (P.B.) for $\alpha=0.1$. Calibration sets all have the size $n=500$.
Comparison with similar methods
We compare our approaches with other methods such as Izbicki et al. [23], Cabezas et al. [6], Martinez et al. [33]. These methods are similar to our approach as they also find data-driven groups on additional datasets for group-conditional conformal prediction. We denote by LOCART and LOFOREST the methods in Cabezas et al. [6], CD-split in Izbicki et al. [23], and TrustyAI Multi Region in Martinez et al. [33]. We use the implementations publically provided by the authors Izbicki et al. [23], Cabezas et al. [6], and Martinez et al. [33]. Note that these methods include their own tuning mechanism/hyperparameter optimization to control the tree structure. For example, LOCART and LOFOREST use cross-validation to tune the complexity parameter used for minimal cost-complexity pruning to minimize the mean-square error in predicting conformity scores. The multi-region method similarly optimizes tree hyperparameters (pruning complexity parameter, minimum number of samples per leaf, etc) via K-fold cross validation to optimize the coverage ratio. The results are in Table 6, which shows that our approach outperforms despite the conceptual similarity. We think that our improvement is attributed to the robust tree, which frees us from any need to split the calibration dataset to find the group information.
Split Conformal Conformal Tree Conformal Forest LOCART LOFOREST CD-split TrustyAI Multi Region
Width Cov. ISL Width Cov. ISL P.B. Width Cov. ISL P.B. Width Cov. ISL P.B. Width Cov. ISL P.B. Width Cov. ISL P.B. Width Cov. ISL P.B.
Data 1 4.16 0.88 6.12 3.27 0.87 4.26 0.6 3.52 0.89 4.11 0.66 3.99 0.88 6.19 0.8 3.99 0.88 6.19 0.8 4.12 0.82 11.1 0.69 3.78 0.85 6.15 1.0
Data 2 2.85 0.9 7.02 3.02 0.86 5.52 0.68 3.87 0.91 4.55 0.75 2.64 0.88 7.06 0.8 2.64 0.88 7.06 0.8 4.5 0.89 12.36 0.47 2.72 0.88 7.08 0.6
bike 1.3 0.9 1.98 1.34 0.88 2.02 0.57 1.11 0.92 1.31 0.76 1.16 0.9 1.48 0.67 1.18 0.91 1.49 0.64 2.25 0.86 4.33 0.08 1.13 0.88 1.58 0.64
bio 1.66 0.9 2.21 1.67 0.9 2.25 0.73 1.51 0.92 1.93 0.56 1.56 0.9 2.1 0.52 1.57 0.91 2.07 0.53 4.07 0.61 0.78 0.78 27.61 1.0 27.61 0.0
community 1.95 0.91 2.81 1.66 0.87 2.6 0.48 1.65 0.9 2.32 0.67 1.74 0.91 2.45 0.67 1.76 0.93 2.43 0.6 2.08 0.85 4.85 0.47 1.99 0.94 2.5 0.56
concrete 0.61 0.91 0.84 0.71 0.89 0.92 0.7 0.77 0.93 0.86 0.54 0.62 0.92 0.82 0.38 0.61 0.91 0.84 0.4 1.9 0.95 2.18 0.0 1.28 0.99 1.31 0.0
homes 0.61 0.9 1.21 0.63 0.89 1.03 0.57 0.62 0.92 0.82 0.72 0.64 0.91 0.93 0.69 0.63 0.92 0.91 0.68 1.26 0.84 3.42 0.0 72.94 1.0 72.94 0.0
star 0.17 0.91 0.21 0.17 0.89 0.22 0.5 0.18 0.9 0.21 0.46 0.17 0.91 0.21 0.47 0.17 0.91 0.21 0.21 0.17 0.85 0.24 0.39 23.02 1.0 23.02 0.0
Table 6: Empirical results on real and synthetic data as the average over five trials on width, coverage, and proportion better (P.B.) for $\alpha=0.1$.
Comparison of prediction set sizes for skin disease prediction
For the skin disease application in Section 5.2, we also draw the frequency of the prediction set size in Figure 9. We can see that the frequency of a single element set is the highest for conformal tree and more concentrated near smaller set sizes, whereas most of predictions of Naive UQ predominate size 5.
Figure 9: Comparison of prediction set sizes for erythemato squamous disease using GPT-4o. Conformal Tree yields prediction sets that are smaller than standard conformal prediction, containing 1.33 less items on average.
Interpretable box-wise visualization
For the smaller tree $(m=50)$ in Section 5.2, we visualize its empirical test coverage, threshold, and prediction confidence in Figure 10. Note that on Box 3, the model prediction quality is high (Figure 10 (i)). On this box, the threshold values (Orange line in Figure 10 (f)) are relatively very small compared to those of split conformal prediction.
Figure 10: Empirical test coverage, threshold, and prediction confidence on each box were obtained by applying Conformal Tree on one realization of the calibration dataset ($60\%$ of the data). The behavior of Conformal Tree in this example is very interpretable: In Box 3, our method has smaller thresholds, improving the coverage rate significantly. In Box 1, the behavior is the opposite. The empirical coverage estimates in (a) are crude due to the lack of samples (see, Figure 11). For the splitting process of this tree, see Figure 7.
Empirical test coverage
In Figure 11, we plot the empirical test coverage of split conformal and Conformal Tree on the GPT diagnosis example in Section 5.2. For a single split (Figure 11 Left), the empirical estimation of the coverages even for the split conformal prediction has a high variation. We think it was because of the number of test data points is small. To alleviate it, we repeat random splits and apply these conformal methods and take average over the estimated coverage for each $\alpha$ value. In Figure 11 Right shows, as the number of random splits increases, the coverage aligns more with the identity line. Note that in any ways, Conformal tree always empirically covers at the desired theoretical coverage level.
@@ Appendix E Additional Experimental Results
Figure 11: Empirical test coverage for split-conformal and Conformal Tree for varying $1-\alpha$ values on the GPT diagnosis example. Left: The empirical coverage rate for a single split. Right: The empirical (average) coverage rate for 40 random splits.
Conformal Tree’s tree structure
In Figure 12, we additionally provide an interpretability visualization of the tree that corresponds to 6. It shows each level of Conformal Tree’s structure on the skin disease calibration data similar to Figure 7.
@@
Figure 12: First five levels of Conformal Tree’s structure on the skin disease calibration data, with $m=10$ and $K_{max}=20$. Node colors correspond to the mid-range of conformity scores within the node. Below each node, we plot the empirical distribution of calibration data corresponding to each disease within that node.